OpenStudy (anonymous):

Solve 3 sin(x/2) - 1 = 0 at interval (0, 2pi)

OpenStudy (anonymous):

\[3 \sin\left(\frac{x}{2} \right) - 1 = 0\]\[3 \sin\left(\frac{x}{2} \right) = 1\]\[\sin\left(\frac{x}{2} \right) = \frac{1}{3}\]

OpenStudy (anonymous):

\[\frac{x}{2} = \sin^{-1} \left(\frac{1}{3} \right)\]\[x = 2 \sin^{-1} \left(\frac{1}{3} \right)\]

OpenStudy (anonymous):

Did I clear everything?

OpenStudy (anonymous):

Yes, thank you but now I havae a cot question

OpenStudy (anonymous):

2 cot^2 x - cot x = 3

OpenStudy (anonymous):

interval (0, 2pi)